F(x, y, z) =< P, Q, R >=<-y +z,x-z,x-y> S: z = 9 - x2 - y2 and z>0 (9a) Evaluate W= $ P dx + Qdy + Rdz с
Answer #1

Sell <-4+2, naz, nagy f(xidize LP, Q, R7 Curl a of is given by - j K Curl(fa day Hay Haz 1-2 ng 1-ytz i(-1+1) -j (1-1) +*(1+1I SS (oto+2) 4A = 2 SS &A 22 + y229 = 2X Area of the f the circle with rad 146 3 A 2 = 22 + x (3) 2 - 2AX 9 I= 187